OpenStudy (anonymous):

find the domain of the function p(x)=x^2-2x+3 What is the domain of P

OpenStudy (anonymous):

\(( \infty, \infty)\).

OpenStudy (anonymous):

\((- \infty, \infty)\)*

OpenStudy (anonymous):

does that mean this is a real number

OpenStudy (anonymous):

Yeah, the domain of \(p(x)\) os the set of all real numbers \( \ R \).

OpenStudy (anonymous):

is*

OpenStudy (anonymous):

so the 3 is a real number as well as 0 meaning x|x is a real number